Na 2 CO 3(s) +2 HCl (aq) 2 NaCl (s) +H 2 O (g) +CO 2(g)

You are working in a laboratory at room temperature , which is considered to be 25 degrees * C The atmospheric pressure of the room from the barometer is 769 mmHg. Calculate the volume in Liters of carbon dioxide produced by the reaction of 12.78g sodium carbonate with 16.8 grams of concentrated (12.1M) hydrochloric acid


1
Expert's answer
2021-04-21T06:09:45-0400

The equation given above is incomplete. The correct one is:

Na2CO3 + 2HCl --> 2NaCl + H2O + CO2

First, the limiting reactant should be determined by comparing the amounts of reactants in moles. The density of concentrated HCl is not given here, but it is estimated to be approximately 1.20 g/mL. So the volume of HCl equals "\frac{16.8g}{1.20g/mL}=14.0mL=0.0140L"


"n(Na_2CO_3)=\frac{12.78g}{105.99g/mol}=0.1206mol"


"n(HCl)=12.1M\times0.0140L=0.169mol"


(0.169 : 0.1206) < (2 : 1), hence HCl is the limiting reactant.


Amount of CO2 in moles:

"n(CO_2)=0.169mol(HCl)\times\frac{1mol(CO_2)}{2mol(HCl)}=0.0845mol"


Unit conversions:

25oC in Kelvins is 25 + 273 = 298 K.

769 mmHg = "\frac{769}{760}=1.012atm"


Finally, solving for V according to the ideal gas law:


"V=\frac{nRT}{P}=\frac{0.0845mol\times0.08206\frac{L\cdot{atm}}{mol\cdot{K}}\times298K}{1.012atm}=2.04L"


Answer: 2.04 L